1. What are the steps to solving this problem ?
2. What would be the correct answer to this problem ?
Problem: (-7)+8(1+2h)=18

Answer by TimothyLamb(4379)

(-7) + 8(1+2h) = 18
---
-7 + 8 + 16h = 18
1 + 16h = 18
16h = 17
Answer:
h = 17/16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php